# ORN-CBF: Learning Observation-conditioned Residual Neural Control Barrier Functions via Hypernetworks

### v2

@@ I INTRODUCTION
CBFs, inspired by control Lyapunov functions, are scalar functions defined over the system’s state space that characterize control invariant sets based on Nagumo’s Theorem on set invariance [2, 3]. In practice, the key advantage of using CBFs is emphasized for control-affine systems, where the safety filtering is reduced to solving a quadratic program (QP) and can be performed with high frequency [4]. However, even though CBFs provide an elegant way to enforce safety, the main difficulty lies in designing a proper CBF. This problem is especially complex for nonlinear systems with state and input constraints, and most of the existing approaches introduce certain assumptions or simplifications to obtain a solution [1]. Also, an additional layer of complexity is present for systems such as mobile robots that operate in unknown environments based on local observations. For such systems, the CBF must be generated in real time using the available observations, meaning that the existing methods for offline CBF design are not applicable [5].
In this paper, we propose a learning-based approach for designing observation-conditioned CBFs that advances beyond the limitations of existing methods thanks to a few key features. First, we use an efficient hypernetwork-based architecture where the hypernetwork parametrizes the main network given available observations, while the main network approximates the target function over the system’s state space. This architecture improves efficiency by inferring the hypernetwork (complex model) only once when a new observation becomes available, while the main network (simple model) is queried with high frequency for its value and gradient. Second, we use Hamilton-Jacobi (HJ) reachability analysis as the source of supervision during training, resulting in CBFs that approximately recover the maximal control invariant set [6, 7, 8]. Finally, instead of directly approximating the HJ value function, the main network approximates only the residual component with respect to the signed distance function (SDF). Since the residual function is always nonnegative [9], by applying a nonnegative activation function at the output of the main network, the safe set associated with the predicted CBF never intersects with the observed failure set. The summary of the contributions is:
@@ II RELATED WORK
There is a significant amount of work on CBFs published in recent years as they have been proven to be a powerful, yet elegant way to enforce safety [3]. Numerous researchers contributed through their work on different variants of CBFs, including exponential CBFs [10], discrete-time CBFs [11], robust CBFs [12], adaptive CBFs [13], etc. Nevertheless, designing a proper CBF for a general nonlinear system with state and control constraints remains the major challenge in the field [1].
The particularly relevant stream of research leverages machine learning to design CBFs in a more flexible and scalable way [14]. For example, learning CBFs from safe trajectories obtained via expert demonstrations is proposed in [15], while the neural CBFs for systems with limited control are introduced in [16]. However, these approaches often fail to recover the maximal safe set, which can be obtained via HJ reachability analysis [8]. The connection between CBFs and HJ reachability is established in [6], where the novel control-barrier value function (CBVF) is introduced as a method for systematic design of CBFs that recover the maximal safe set. This approach is further extended in [7] to iteratively refine candidate CBFs, hand-crafted or learned, in order to enlarge the initial safe set. Similarly, authors in [17] use HJ reachability to locally modify potentially unsafe regions associated with a neural CBF, improving the overall scalability. Still, these methods have been demonstrated only for systems operating in known environments, which is the key difference to our method.
Regarding the design of neural CBFs in unknown environments, the method in [18] learns a neural approximation of an SDF for the given observation and uses it as a CBF. The main drawback is that the training is performed online, limiting its application to systems with slower dynamics. Differently, the authors in [19] use a discrete-time CBF, which requires predicting the sensor observation one time step into the future. This approach is feasible for simple sensors such as 2D laser scan, but might be impractical for more complex modalities. On the other hand, authors in [5] train a neural model to approximate the value function associated with the state-dependent Riccati equation based on a 2D laser scan and use it as CBF.
@@ III-B Control Barrier Functions
which is linear in $u$ for a given $x$. Moreover, if the set of admissible controls $\mathcal{U}$ can be represented by a set of linear constraints, the CBF condition 4 can be used to design a safety filter in the form of a quadratic program (CBF-QP) [1, 4]. The CBF-QP safety filter provides safe control $u^{*}$ by solving the following optimization problem for the nominal control $u$ and state $x$:
$u^{*}=\operatorname*{arg\,min}_{u^{*}\in\mathcal{U}}$ $\frac{1}{2}\|u-u^{*}\|^{2}$ (5)
@@ IV-B Observation-conditioned residual neural CBF
We approximate the residual using a neural network with parameters $\Theta$ and denote it with $\hat{r}_{\Theta}(x|o)$. To ensure a nonnegative value for all inputs, we apply a nonnegative activation at the model’s output. A straightforward option would be the ReLU function defined as $\max(0,z)$. However, the gradient of the function is always equal to 0 for negative inputs, which can negatively impact the training process. Therefore, we opt for a smooth approximation of the ReLU function and in this paper we use the softplus function:
@@
In addition, given that the approximation of the CBF gradient $\nabla\hat{h}(x|o)$, and therefore $\nabla\hat{r}_{\Theta}(x|o)$, is required for the CBF constraint 13, we use the multilayer perception (MLP) model with sinusoidal activation functions. Those models have been proven to work well for applications where the gradient of the approximated function has to be accurately approximated as well [22, 23].
@@ V-A Robot Models
with linear speed $v=1.0$ m/s and control limit $\omega\in[-0.5,0.5]$} rad/s.
@@ V-B Model Training
As described in Section IV, we train the model in a supervised fashion, and the first part is to obtain a suitable dataset consisting of inputs and target outputs. For the hypernetwork inputs, we first collect a set of observations in the form of occupancy grid maps and compute the corresponding discretized SDFs. For the ground robot, we collect 2,500 costmaps of size $\times6$ m with resolution 0.06 m from a 3D model of a warehouse environment. Similarly, we collect 2,000 costmaps of size $\times6$ m with resolution 0.075 m from a forest-like 3D environment for the quadcopter model. We additionally augment both datasets by flipping horizontally and rotating for 90, 180 and 270 each original and flipped costmap, resulting in 20,000 costmaps for the ground robot and 16,000 costmaps for the quadcopter.
@@
Next, we define state space grids over which the HJ value functions are computed and which are used as the input to the main network. We define state space grids of shape $100\times 100\times 30$} for the Dubins car and $80\times 80\times 20\times 20$} for the 2D double integrator model. To compute the HJ value functions, we use the hj_reachability [^5] https://github.com/StanfordASL/hj_reachability Python library. After the computation is performed, we store the value functions and grid points only for the subset of the initial grid according to the explanation in the last paragraph of Section IV. Concretely, for both models, we store only the central patch in the positional subspace of size $16\times 16$}, meaning that we store $16\times 16\times 30$} grid points and the corresponding function values for the Dubins car and $16\times 16\times 20\times 20$} for the 2D double integrator.
@@
After aggregating the datasets, we proceed with the model training. For both types of robots, we use deep CNN models as the hypernetwork and MLP models as the main network. We use the same architecture of the main network in both cases: 32-32-32-16-16-16-8-8-8 units with sinusoidal activations in hidden layers and a single output with softplus activation, while the number of inputs is equal to the corresponding state dimension. During the training, at each epoch, a discretized SDF is propagated through the hypernetwork and the main network is parametrized with the predicted parameters. Then, the complete state space grid is propagated through the main network to obtain the HJ value function predictions. Finally, the loss function is evaluated for the predicted and true value functions, and the parameters of the hypernetwork are updated via the gradient descent method. In this paper, we use the RWMSE loss function proposed in [20] and train models for 100 epochs with a batch size of 16. We use the Adam optimizer with an initial learning rate of $10^{-4}$, which is decreased by a factor of 10 at epochs 85 and 95. The training process takes $\sim$0.5 h for the Dubins car and $\sim$2.9 h for the 2D double integrator on a single NVIDIA RTX 3090 GPU, which is roughly 20 times faster compared to the method in [20] under the same setting. In Fig. 4 we visualize a sample input to the hypernetwork and the corresponding learned ORN-CBF for the Dubins car model.
Figure 4: An example of a learned ORN-CBF for the Dubins car model.
@@ V-C Simulation Experiments
The proposed ORN-CBF method is first evaluated in simulations implemented using the Gazebo 3D simulator. The ground robot, modeled as a Dubins car, is tested in a warehouse environment visualized in Fig. 5. As a nominal controller, we use a simple MPC local planner, which uses the SDF as a collision avoidance constraint (SDF-MPC). The ORN-CBF method is examined against SDF-MPC alone and more advanced MPC-based methods that have been proven to work well in practice for safe navigation in unknown environments: DCBF-MPC [24] and NTC-MPC [20]. The related works on neural CBFs [19] and [5] assume different observation modalities, and therefore are not directly applicable in our setting. Besides that, as an ablation study, we test a simpler variant of the proposed method, which directly approximates the HJ value function by the main network instead of the residual function. We call this method observation-conditioned neural CBF (ON-CBF). Even though this version has a less complex architecture, it does not provide any safety guarantees on the approximation error. As it is standard in practice, we use a linear function for $\alpha(\cdot)$, i.e. $\alpha(h)=kh,k\in\left[0,\infty\right)$.
@@ V-D Hardware Experiments
We perform hardware experiments with a ground robot modeled as a Dubins car, which is shown in Fig. 10. We use exactly the same ORN-CBF and ON-CBF models trained on the synthetic data to demonstrate the robustness of the proposed method when transferring from simulation to the real world. We compare it against the same MPC-based baselines with the prediction horizon of 10 steps, sampling time of 0.1s, and control frequency of 20 Hz. The CBF safety filters run at 200 Hz and the complete computation is performed onboard. We conduct 10 experiments with random distributions of obstacles in the environment for each method. The obtained results are presented in Table I, demonstrating significantly better results of the proposed safety filters compared to the baselines. Fig. 11 shows the change of CBF value over time for ORN-CBF and ON-CBF methods for one scenario from the experiments.
@@ VI DISCUSSION ON LIMITATIONS
For the proposed method, we use numerical tools for HJ reachability analysis during the data generation process. Due to the curse of dimensionality, those tools are usually not applicable for systems with more than six state dimensions. To improve scalability, one can use the self-supervised approach for model training as in [23], or compute approximate HJ value functions using the MPC framework [26].
@@ VII CONCLUSIONS
We propose a novel learning-based method for designing observation-conditioned CBFs in unknown environments. The resulting CBF approximately recovers the maximal safe set corresponding to the current observation. By learning only the residual component of the HJ value function, it is guaranteed that the predicted safe set never includes the observed failure set. Moreover, the proposed hypernetwork-based architecture allows for a computationally efficient implementation of the safety filter.
The proposed ORN-CBF method and its simpler variant, ON-CBF, are extensively evaluated in simulations for a ground robot and a quadcopter. The results indicate improved success rates and robust performance in out-of-domain environments compared to the baselines. The hardware experiments demonstrate that our safety filters can be easily deployed on different real robots. In future work, we plan to extend the proposed method to dynamic environments and higher-dimensional systems.

### v3

@@ I INTRODUCTION
CBFs, inspired by control Lyapunov functions, are scalar functions defined over the system’s state space that characterize control invariant sets based on Nagumo’s Theorem on set invariance [2, 3]. In practice, the key advantage of using CBFs is emphasized for control-affine systems, where the safety filtering is reduced to solving a quadratic program (QP) and can be performed with high frequency [4]. Although CBFs provide an elegant way to enforce safety, the main difficulty lies in designing a proper CBF. This problem is especially complex for nonlinear systems with state and input constraints, and most existing approaches make assumptions or simplifications to obtain a solution [1]. Also, an additional layer of complexity is present for systems such as mobile robots that operate in unknown environments based on local observations. For such systems, the CBF must be generated in real time using the available observations, meaning that the existing methods for offline CBF design are not applicable [5].
In this paper, we propose a learning-based approach for designing observation-conditioned CBFs that overcome the limitations of existing methods by leveraging a few key features. First, we use an efficient hypernetwork-based architecture in which the hypernetwork parametrizes the main network given available observations (2D occupancy grid in our experiments), while the main network approximates the target function over the system’s state space. This architecture improves efficiency by inferring the hypernetwork (complex model) only once when a new observation becomes available, while the main network (simple model) is queried frequently for its value and gradient. Second, we use Hamilton-Jacobi (HJ) reachability analysis as the source of supervision during training, resulting in CBFs that approximately recover the maximal control invariant set [6, 7, 8]. Finally, instead of directly approximating the HJ value function, the main network approximates only the residual component with respect to the signed distance function (SDF). Since the residual function is always nonnegative [9], by applying a nonnegative activation function at the output of the main network, the safe set associated with the predicted CBF never intersects with the observed failure set. The summary of the contributions is:
@@ II RELATED WORK
There is a significant amount of work on CBFs published in recent years as they have been proven to be a powerful, yet elegant way to enforce safety [3]. Numerous researchers contributed through their work on different variants of CBFs, including exponential, discrete-time, robust, time-varying, and adaptive CBFs [10]. Nevertheless, designing a proper CBF for a general nonlinear system with state and control constraints remains the major challenge in the field [1].
The particularly relevant stream of research leverages machine learning (ML) to design CBFs in a more flexible and scalable way. For example, supervised ML is applied in [11] to construct CBFs from LiDAR rays, learning CBFs from expert trajectories is proposed in [12], while the neural CBFs for systems with limited control are introduced in [13]. Also, an uncertainty-aware CBF method based on Gaussian processes is developed in [14]. However, these approaches often fail to recover the maximal safe set, which can be obtained via HJ reachability analysis [8]. The connection between CBFs and HJ reachability is established in [6], where the novel control-barrier value function (CBVF) is introduced. This approach is further extended in [7] to iteratively refine candidate CBFs, hand-crafted or learned, to enlarge the initial safe set. Similarly, authors in [15] use HJ reachability to locally modify potentially unsafe regions associated with a neural CBF, improving the overall scalability. Still, these methods have been demonstrated only for systems operating in known environments, which is the key difference to our method.
Regarding the design of neural CBFs in unknown environments, the method in [16] learns a neural SDF approximation for the given observation and uses it as a CBF. The main drawback is that the training is performed online, limiting its application to systems with slower dynamics. Differently, the authors in [17] use a discrete-time CBF, which requires predicting the sensor observation one time step into the future. This approach is feasible for simple sensors such as 2D laser scan, but might be impractical for more complex modalities. On the other hand, authors in [5] train a neural model to approximate the value function associated with the state-dependent Riccati equation based on a 2D laser scan and use it as CBF.
@@ III-B Control Barrier Functions
which is linear in $u$ for a given $x$. Moreover, if the set of admissible controls $\mathcal{U}$ can be represented by a set of linear constraints, the CBF condition 4 can be used to design a safety filter in the form of a quadratic program (CBF-QP) [1, 4]. The CBF-QP safety filter provides safe control $u^{*}$ by solving the following optimization problem for the nominal control $u_{ref}$ and state $x$:
$u^{*}=\operatorname*{arg\,min}_{u\in\mathcal{U}}$ $\frac{1}{2}\|u_{ref}-u\|^{2}$ (5)
@@ IV-B Observation-conditioned residual neural CBF
We approximate the residual using a neural network with parameters $\Theta$ and denote it with $\hat{r}_{\Theta}(x|o)$. To ensure a nonnegative value for all inputs, we apply a nonnegative activation at the model’s output. A straightforward option would be the ReLU function defined as $\max(0,z)$. However, the gradient of the function is always equal to 0 for negative inputs, which can negatively impact the training process. Also, the activation should be continuously differentiable, so we opt for the softplus function:
@@
In addition, given that the approximation of the CBF gradient $\nabla\hat{h}(x|o)$, and therefore $\nabla\hat{r}_{\Theta}(x|o)$, is required for the CBF constraint 13, we use the multilayer perceptron (MLP) model with sinusoidal activation functions. Those models have been proven to work well for applications where the gradient of the approximated function has to be accurately approximated as well [20, 21].
@@ V-A Robot Models
where $v=1.0$ m/s and $\omega\in[-0.5,0.5]$} rad/s.
@@ V-B Model Training
As described in Section IV, we train the model in a supervised fashion, and the first part is to obtain a suitable dataset consisting of inputs and target outputs. For the hypernetwork inputs, we first collect a set of observations in the form of occupancy grid maps and compute the corresponding discretized SDFs using the method described in [22]. For the ground robot, we collect 2,500 costmaps of size $\times6$ m with resolution 0.06 m from a 3D model of a warehouse environment. Similarly, we collect 2,000 costmaps of size $\times6$ m with resolution 0.075 m from a forest-like 3D environment for the quadcopter model. We additionally augment both datasets by flipping horizontally and rotating for 90, 180 and 270 each original and flipped costmap, resulting in 20,000 costmaps for the ground robot and 16,000 costmaps for the quadcopter.
@@
Next, we define state space grids over which the HJ value functions are computed and which are used as the input to the main network. We define state space grids of shape $100\times 100\times 30$} for the Dubins car and $80\times 80\times 20\times 20$} for the 2D double integrator model. To compute the HJ value functions, we use the hj_reachability [^5] https://github.com/StanfordASL/hj_reachability Python package. The computation takes $\sim$1.17 h in the first case, and $\sim$16.3 h in the second case, on a single RTX3090 GPU. Then, we store the value functions and grid points only for the subset of the initial grid according to the explanation in the last paragraph of Section IV. Concretely, for both models, we store only the central patch in the positional subspace of size $16\times 16$}, meaning that we store $16\times 16\times 30$} grid points and the corresponding function values for the Dubins car and $16\times 16\times 20\times 20$} for the 2D double integrator.
@@
After aggregating the datasets, we proceed with the model training. For both types of robots, we use deep CNN models as the hypernetwork and MLP models as the main network. We use the same architecture of the main network in both cases: 32-32-32-16-16-16-8-8-8 units with sinusoidal activations in hidden layers and a single output with softplus activation, while the number of inputs is equal to the corresponding state dimension. During the training, at each epoch, a discretized SDF is propagated through the hypernetwork and the main network is parametrized with the predicted parameters. Then, the cropped state space grid is propagated through the main network to obtain the HJ value function predictions. Finally, the loss function is evaluated for the predicted and true value functions, and the parameters of the hypernetwork are updated via the gradient descent method. In this paper, we use the RWMSE loss function proposed in [18] and train models for 100 epochs with a batch size of 16. We use the Adam optimizer with an initial learning rate of $10^{-4}$, which is decreased by a factor of 10 at epochs 85 and 95. The training process takes $\sim$0.5 h for the Dubins car and $\sim$2.9 h for the 2D double integrator on a single RTX3090 GPU, which is roughly 20 times faster compared to the method in [18] under the same setting. In Fig. 4 we visualize a sample input to the hypernetwork and the corresponding learned ORN-CBF for the Dubins car model.
Figure 4: A slice of a learned ORN-CBF for the Dubins car for $\theta=\frac{\pi}{2}$ rad.
@@ V-C Simulation Experiments
The proposed ORN-CBF method is first evaluated in simulations implemented using the Gazebo 3D simulator. The ground robot, modeled as a Dubins car, is tested in a warehouse environment visualized in Fig. 5. As a nominal controller, we use a simple MPC local planner, which uses the SDF as a collision avoidance constraint (SDF-MPC). The ORN-CBF method is examined against SDF-MPC alone and more advanced MPC-based methods that have been proven to work well in practice for safe navigation in unknown environments: DCBF-MPC [23] and NTC-MPC [18]. The related works on neural CBFs [17] and [5] assume different observation modalities, and therefore are not directly applicable in our setting. Besides that, as an ablation study, we test a simpler variant of the proposed method, which directly approximates the HJ value function by the main network instead of the residual function. We call this method observation-conditioned neural CBF (ON-CBF). Even though this version has a less complex architecture, it does not provide any guarantees on the approximation error. As it is standard in practice, we use a linear function for $\alpha(\cdot)$, i.e. $\alpha(h)=kh,k\in\left[0,\infty\right)$.
@@ V-D Hardware Experiments
We perform hardware experiments with a ground robot modeled as a Dubins car, shown in Fig. 10. To evaluate sim-to-real robustness, we use the same ORN-CBF and ON-CBF models trained on synthetic data. We compare against the same MPC-based baselines using a prediction horizon of 10 steps, a sampling time of 0.1 s, and a control frequency of 20 Hz. The CBF safety filters run at 200 Hz, and all computations are performed onboard. For each method, we run 10 experiments with random obstacle distributions. The results in Table I show that the proposed safety filters perform significantly better than the baselines. Fig. 11 illustrates the CBF value over time for ORN-CBF and ON-CBF in one representative scenario. The CBF values occasionally dip slightly below zero due to model mismatch and measurement noise, yet collisions are avoided thanks to a small buffer zone that accounts for these uncertainties.
@@ VI DISCUSSION ON LIMITATIONS
We use numerical tools for HJ reachability during data generation, but these tools are typically not applicable to systems with more than six state dimensions. To improve scalability, one can use the self-supervised approach for model training as in [21], or compute approximate HJ value functions using the MPC framework [26].
@@ VII CONCLUSIONS
We propose a novel learning-based method, called ORN-CBF, for designing observation-conditioned CBFs in unknown environments. The resulting CBF represents an HJ value function and approximately recovers the maximal safe set for the current observation. By learning only the residual component of the HJ value function, we guarantee that the predicted safe set never includes the observed failure set. Moreover, the proposed hypernetwork-based architecture enables a computationally efficient safety filter.
We extensively evaluate ORN-CBF in simulations on a ground robot and a quadcopter. The results show improved success rates and robust out-of-domain performance compared to the baselines. Hardware experiments further demonstrate that the safety filters can be readily deployed on different real robots. Future work will extend the method to dynamic environments and higher-dimensional systems.